How do i solve a problem like this?:

Evaluate x² - 3x + 4 evaluated for x=1 + i equal

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

Evaluate x^2-3x+4 evaluated for x=1+i
-----------
Sub 1+i for x
x^2-3x+4
= (1+i)^2 - 3(1+i) + 4
= 1 + 2i -1 - 3 - 3i + 4
= -i + 1
= 1 + i

Answer by Edwin McCravy(20055)

Evaluate x² - 3x + 4 evaluated for x=1 + i


                  x² - 3x + 4

Replace each x by (1 + i):

      (1 + i)² - 3(1 + i) + 4

  (1 + i)(1 + i) - 3 - 3i + 4

(1 + i + i + i²) - 3i + 1

   (1 + 2i + i²) - 3i + 1
 
     1 + 2i + i² - 3i + 1

      2 - i + i²

Then you replace i² by (-1)

      2 - i + (-1)

        2 - i - 1
               
            1 - i

Edwin